a charge of 0.20 microC is 30 cm from a point charge of 3.0 microC in vacuum. what work is required to bring the 0.20microC charge 18 cm closer to the 3.0microC charge?

the answer must be 0.027 J
Answer #1

here,

the charge 1 , q1 = 0.2 uC = 0.2 * 10^-6 C

the charge 2 , q2 = 3 uC = 3 * 10^-6 C

initial seperation , r0 = 30 cm = 0.3 m

final seperation , r = 18 cm = 0.18 m

the work done required , W = change in potential energy

W = K * q1 * q2 /(r0 - r) - K * q1 * q2 /r0

W = 9 * 10^9 * 0.2 * 10^-6 * 3 * 10^-6 * ( 1/(0.3 - 0.18) - 1/0.3) J

W = 0.027 J